you roll a dice twice. What is the probability of getting pairs of faces such that the sum of the two faces is a prime number?
A 15/36 B 16/36 C 17/36 D 18/36 E 19/36



Answer by Fombitz(32388)

There are 36 possible outcomes from rolling two dice.
You can roll any number from 2 to 12 or 11 possible outcomes.
5 of the 11 are prime (2,3,5,7,11).
Rolling 2: (1,1): 1 outcomes
Rolling 3: (1,2)(2,1): 2 outcomes
Rolling 5: (1,4)(2,3)(3,2)(4,1): 4 outcomes
Rolling 7: (1,6)(2,5)(3,4)(4,3)(5,2)(6,1): 6 outcomes
Rolling 11: (5,6)(6,5): 2 outcomes
Total number of prime outcomes=1+2+4+6+2=14
.
.
.
P(prime sum)=15/36=5/12
.
.
.
A) is the correct answer.